OpenStudy (thecuteone25):

Carolyn drove 150 miles in 3 hours. Which rate represents this situation?

OpenStudy (thecuteone25):

I will Medal

OpenStudy (yoloshroom):

Distance = Rate x Time?

OpenStudy (yoloshroom):

150 = R x 3 150 = 3R 150/3 = R Not sure if that's all u needed, or if im off :)

OpenStudy (thecuteone25):

That helped Thank you

OpenStudy (michele_laino):

the requested rate is given by the subsequent computation: \[\huge r = \frac{{{\text{distance}}}}{{{\text{time}}}} = \frac{{150}}{3} = ...?\]

OpenStudy (thecuteone25):

Who ever tries to help I will medal or even for being on my question